OpenStudy (anonymous):

write the equation of a line with the given properties in slope intercept form slope= -3 through (-4,9)

OpenStudy (anonymous):

Start by writing it in point-slope form: (Y-Y1)=M(X-X1) In this instance that would be (Y-9)=-3(X+4) Solve for Y: Y=-3x-3

OpenStudy (anonymous):

thanks